# Privacy-Aware Semantic Cache for Large Language Models

### v1

@@ Abstract.
Large Language Models (LLMs) like ChatGPT, Google Bard, Claude, and Llama 2 have revolutionized natural language processing and search engine dynamics. However, these models incur exceptionally high computational costs. For instance, GPT-3 consists of 175 billion parameters and inference on these models also demands billions of floating-point operations. Caching is a natural solution to reduce LLM inference costs on repeated queries. However, existing caching methods are incapable of finding semantic similarities among LLM queries, leading to unacceptable false hit-and-miss rates.
This paper introduces MeanCache, a semantic cache for LLMs that identifies semantically similar queries to determine cache hit or miss. Using MeanCache, the response to a user’s semantically similar query can be retrieved from a local cache rather than re-querying the LLM, thus reducing costs, service provider load, and environmental impact. Existing LLM Cache trains a single model at the server to identify similarities across multiple queries, thus suffering from privacy, cache scalability, and wasteful query requests. MeanCache leverages Federated Learning (FL) to collaboratively train a query similarity model in a distributed manner across numerous users without violating privacy. By placing a local cache in each user’s device and using FL, MeanCache reduces the latency and costs and enhances model performance, resulting in lower cache false hit rates. MeanCache compresses the embedding dimensions to minimize cache storage and also finds the optimal cosine similarity threshold. Our experiments, benchmarked against the GPTCache, reveal that MeanCache attains an approximately 17% higher F-score and a 20% increase in precision during semantic cache hit-and-miss decisions. Furthermore, MeanCache reduces the storage requirement by 83% and accelerates semantic cache hit-and-miss decisions by 11%, while still surpassing GPTCache.
@@ 1. Introduction
Large Language Models (LLMs) like ChatGPT (cha, [n.d.]), Google Bard (bar, [n.d.]), Claude (cla, [n.d.]), and Llama 2 (Touvron et al., 2023) have demonstrated remarkable capabilities in understanding and generating human language, leading to significant advancements in a variety of applications ranging from search engines to conversational agents. The integration of LLMs into platforms like the Perplexity AI search engine  (per, [n.d.]b), Rabbit OS (rab, [n.d.]), Arc browser (arc, [n.d.]), and Bing Chat (bin, [n.d.]) has underscored their growing influence. For instance, in 2023, 1.9 billion queries were sent to Bing chat, reflecting the massive use of such systems (bin, [n.d.]).
@@
Motivation. Traditional search engines respond to user queries with ranked result pages based on keyword matching (Lempel and Moran, 2003; Jansen et al., 2000; Markatos, 2001). In contrast, LLM-powered services like Perplexity AI search engine, Bing chat, and ChatGPT provide a single precise answer to a user query. However, generating these responses with LLMs, such as GPT-3, requires substantial computations and poses environmental challenges (Frantar et al., 2023; Wang et al., 2023; Tseng et al., 2024; Sachdeva et al., 2024; Ma et al., 2024; Strubell et al., 2019). For example, GPT-3’s 175B parameters in float16 format consume 326 GB memory, exceeding single GPU capacities and necessitating multi-GPU setups (Frantar et al., 2023). These requirements lead to high operational costs. Google Bard’s costs, for instance, can be ten times those of standard keyword searches, potentially adding an extra \$6 billion for half of Google’s search queries (AIi, [n.d.]). Consequently, LLM-based services charge users and limit query rates (per, [n.d.]a; ope, [n.d.]).
@@
Users frequently submit repeated queries to web services (Lempel and Moran, 2003; Xie and O’Hallaron, 2002), with approximately 33% of search engine queries being resubmitted (Markatos, 2001). Caching serves as an effective technique in web services to address these duplicate queries, avoiding redundant computations, significantly improving response time, reducing the load on query processors, and enhancing network bandwidth utilization (Markatos, 2001; Saraiva et al., 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003; Brin and Page, 1998; Podlipnig and Böszörmenyi, 2003; Baeza-Yates et al., 2007). Applying caching to serve repeated queries in LLMs based web services can have a substantial impact, reducing billions of floating point operations, thereby decreasing operational costs and environmental impacts.
@@
Problem. Existing keyword matching caching techniques  (Saraiva et al., 2001; Markatos, 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003) will fail to capture the semantic similarity among repeated queries to LLM based web services and will result in a significantly low hit rate. For instance, if a user sends the query ”How can I increase the battery life of my smartphone?” and later sends a similar query, ”Tips for extending the duration of my phone’s power source.”, the existing caches will not detect the semantic similarity between these two queries and will lead to a cache miss.
Zhu et al. (2023) and GPTCache(Bang, 2023) present server-side semantic caching for the LLMs based services to address the limitations of keyword-matching caching techniques. They determine if a new query is semantically similar to any query in the cache by using an embedding model. For instance, as per study in (Zhu et al., 2023), if a similar query is found, the server returns the response from the cache. Otherwise, a model multiplexer is used to select the most suitable LLM for the query to generate the response.
@@
These semantic caches have several limitations. (1) They demand a significantly large central cache to store the queries and responses of all users, which violates users privacy. (2) They still incur the network cost of sending a user query to the server even if there is a cache hit. (3) The server-side embedding model to find the semantic similarity among queries is generic and is not tailored to each user’s querying patterns (i.e., model does not adapt to the behavior of each user). (4) Lastly, an end user will still be charged for the query even if the query is served from the server cache without sending it to the LLM.
@@
Key Insights and Contributions of MeanCache. Our work, MeanCache, introduces a novel user-centric semantic caching system that tackles various limitations and challenges inherent in semantic search caches today. An effective semantic search cache must be efficient, compact, and privacy-preserving. MeanCache provides a privacy-preserving caching system that can return the response to repeated queries directly from the local cache of the user, bypassing the need to query the LLM based web service. MeanCache’s achieves these goals in the following ways:
The embedding model used to compute a query’s embedding vector should be compact and efficient, enabling swift computation of embeddings. MeanCache achieves this by employing a smaller transformer model, MPNet (Song et al., 2020), to efficiently compute the user’s query embeddings. MPNet’s smaller size and faster operation, compared to larger models like Llama 2 (Touvron et al., 2023), make it compatible with resource-limited user devices, thereby enabling a resource-efficient semantic cache. Our insight is that a smaller model can achieve performance comparable to larger models on custom tasks  (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). To customize the embedding model, MeanCache uses federated learning for training on user data without storing the data on the service provider. This privacy-preserving training not only customizes the embedding model to the user’s querying patterns but also enhances the performance (i.e., accuracy) of semantic caching without compromising user privacy (i.e., without storing user-data at the web server). The runtime performance of MeanCache is primarily influenced by the time taken to match a new query embedding vector with existing ones in the cache to find a semantically similar query. The search time is directly proportional to the dimensions of the embedding vector. To optimize runtime performance, MeanCache compresses the embedding vector by leveraging principal component analysis (PCA) (Pearson, 1901; Hotelling, 1933; Jolliffe and Cadima, 2016), effectively reducing the size of the embedding vector (i.e., projecting it to lower dimensional space).
In MeanCache, embeddings are computed for new queries and their similarity to cached queries is measured using cosine similarity, an effective metric for capturing semantic similarity between texts. However, determining an optimal threshold for cosine similarity is crucial for accurately identifying true cache hits and misses. For example, a low threshold value may result in a high false hit rate, incorrectly matching a new query with cached ones. MeanCache mitigates this by identifying the optimal threshold from the user’s cached queries, specifically the one that maximizes the F score, thereby minimizing the false hit-and-miss rates on the client’s local data. The cost of finding this optimal threshold is approximately equivalent to one epoch of training on the client’s local data. It’s important to note that the optimal threshold, along with the local weights of the client’s embedding model (post local training), is sent to the central server to update the federated learning global model and the global cosine similarity threshold. This global threshold proves beneficial for similarity matching for new users who do not have cached queries. When a new query matches a cached one on a given cosine similarity threshold, MeanCache results in cache hits and delivers responses from the local cache. If not, it processes cache misses by querying the service provider and storing new query-response pairs for future use.
Evaluations. We compare MeanCache with GPTCache(Bang, 2023), a widely-used open-source semantic cache for LLM-based web services. GPTCache (Bang, 2023) is the most closely related work to MeanCache and has garnered over 6,000 stars on GitHub(zil, [n.d.]). We benchmark MeanCache’s performance against GPTCache on the GPTCache’s dataset (dat, [n.d.]) to demonstrate its effectiveness and highlight the improvements MeanCache offers over existing solutions. Our experiments utilize the Flower FL framework (Beutel et al., 2020) and the SBERT library(Reimers and Gurevych, 2019) to train embedding models in a distributed and privacy-preserving manner with 20 clients. The client size is constrained by the limited size of the GPTCache dataset, which is insufficient to distribute among hundreds of clients. However, our findings are not dependent on the number of clients, and MeanCache’s evaluation setup aligns with the prevailing evaluation standard for new studies in FL (Avdiukhin and Kasiviswanathan, 2021; Wang et al., 2020a). MeanCache surpasses GPTCache (Bang, 2023) by achieving a 17% higher F score and approximately a 20% increase in precision in end-to-end deployment for identifying duplicate queries to LLM-based web services. MeanCache’s embedding compression utility approximately reduces storage and memory needs by 83% and results in 11% faster semantic matching, while still outperforming the state-of-the-art GPTCache.
In summary, MeanCache introduces the design of the first user-centric semantic cache that utilizes federated learning. This cache serves repeated queries to LLM-based web services directly from a user’s local cache, thereby eliminating the need to query web services. Users’ queries are not shared with service providers or stored on centralized servers. MeanCache effectively enhances the embedding model to generate high-quality query embeddings using federated learning. This improvement not only boosts MeanCache’s accuracy, exemplified by an increased true hit rate, but also conserves the user’s query quota and reduces the redundant computations required by LLM services, all without compromising user privacy.
@@ 2.1. Federated Learning
Federated Learning is a distributed training of a machine learning (ML) model. The key benefit of this approach is that users’ data never leaves its source. It mainly consists of four steps. In the first step, the central server shares the same model with all the participating clients. In the second step, each client trains the model on its local data. In the third step, the clients send the updated model to the central server. In the fourth step, the central server aggregates the received models’ weights to form a new global model. The aggregation is done mainly by averaging the weights of the models, and there are different aggregating/fusion algorithms, such as FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), and FedMA (Wang et al., 2020b). The FedAvg algorithm is the most popular one, and it aggregates the models with the following equation:
@@
In the above equation, $W_{global}^{t+1}$represents the new global model, $w_{k,t}$ denotes the model of the $k^{th}$ client at the $t^{th}$ round, $n_{k}$ are the number of samples of the $k^{th}$ client, and $n$ is the total number of samples of clients participating in the given round. Steps 1-4 constitute a single Federated Learning training round. This procedure is reiterated for multiple rounds until the model reaches convergence.
@@ 2.2. Transformer Architecture
The transformer architecture is a deep learning model based on the attention mechanism, which allows the model to focus on the important parts of the input (Vaswani et al., 2017). The transformer architecture is primarily used for sequence-to-sequence tasks in NLP, such as machine translation, text summarization, and question-answering. It consists of two main components: the encoder and the decoder. The encoder takes the input sequence and encodes it into a vector representation. The decoder takes the encoded vector representation and decodes it into the output sequence. Both the encoder and the decoder are composed of multiple layers. Each layer in the transformer architecture has two sub-layers: the multi-head attention layer and the feed-forward layer. The multi-head attention layer performs the attention mechanism, which allows the model to focus on different parts of the input sequence. The attention mechanism calculates the importance of each input element by considering its relationship with other elements in the sequence. This enables the model to capture long-range dependencies and improve performance on tasks that require an understanding of context. The feed-forward layer is a fully connected layer that applies a non-linear transformation to the output of the multi-head attention layer. It helps the model capture complex patterns and relationships in the data. The transformer architecture also incorporates positional encoding, which provides information about the position of each element in the input sequence. This helps the model understand the order of the elements, which is crucial for tasks like machine translation.
@@ 2.2.1. Embeddings and Cosine Similarity
3. MeanCache-A User-Centric Semantic Caching System for LLMS
Caching is a well-known technique for improving the performance of content serving systems (Markatos, 2001; Saraiva et al., 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003). Caching at the end-user level can save computation time, and similar repeated queries can be served from the local cache. However, designing a user-centric semantic caching system for the users who engage with LLM-assisted web services, such as Perplexity AI, Bing and ChatGPT, is not trivial and poses several challenges and many factors to consider. Following are the challenges and factors to consider while designing a caching system for LLMs.
3.1. Research Challenges
Web server caching techniques (Baeza-Yates and Saint-Jean, 2003; Long and Suel, 2005; Saraiva et al., 2001; Lempel and Moran, 2003; Marin et al., 2010) primarily utilize keyword matching, returning a list of links that align with the user’s query. However, these techniques fall short when compared to Language Learning Model (LLM) based systems like ChatGPT or the Perplexity AI search engine, which provide a unified response. This discrepancy becomes evident when a user sends similar but not identical queries. For example, queries like ”Draw a line-plot in Python?” and ”Plot a linear graph using Python” may yield suboptimal results, leading to false misses.
GPTCache(Bang, 2023) and Zhu et al. (2023) propose server-side semantic caching for LLM-based services to return a response from the cache if the query is semantically similar to any query in the cache. This approach reduces the computational cost of LLMs. However, these techniques primarily designed for server-side and store users’ queries on the server which is privacy risk. Furthermore, a user will still be charged for the query even if the query is served from the cache. Additionally, GPTCache(Bang, 2023) also propose Llama 2 to generate superior embeddings, thereby improving semantic matching accuracy. However, this proposal has several limitations. Large language models like Llama 2 are not only sizeable, being gigabytes in size, but they also demand substantial computational resources to generate embeddings. The deployment of such models is impractical due to their size and significant computational overhead for semantic matching. Moreover, the large dimension of Llama 2’s embeddings (Llama 2 embeddings dimension is 4096) can cause considerable overhead when matching a new user query embeddings with cached queries embeddings, as the search time is proportional to the dimensions of the embedding vector. Additionally, high-dimensional embeddings also utilize more memory and storage (e.g., the embeddings generated by Llama 2 for a single query requires approximately 32.05 KBs). An alternative could be to use a smaller pre-trained embedding model on the user side. However, these models may not offer the same level of accuracy and are unable to utilize the vast amount of distributed data available on the user’s device. Storing user queries at the central server and fine-tuning smaller embedding model raises privacy concerns as the user queries are stored at the central server.
Even having optimal embedding model, the task of determining whether a new query embeddings are similar to locally cached queries embeddings is challenging. After generating query embeddings using a transformer model, a similarity metric such as cosine similarity is used to ascertain if the new query embeddings match the cached embeddings of previous queries. This process involves setting a threshold for cosine similarity, which is a delicate balance. A low threshold may result in false positives, matching queries to irrelevant cached queries, while a high threshold could lead to false negatives, missing too many real matches. Therefore, setting an appropriate threshold is both mandatory and challenging.
In conclusion, an efficient user-centric semantic cache system should operate on the user side, utilizing a transformer embeddings model with low inference time to generate embeddings. It should autonomously determine the optimal threshold for cosine similarity to enhance the accuracy of embeddings matching to find similar queries. Additionally, the semantic search time of finding similar queries from the cache should be optimized. Ultimately, this caching system should leverage the vast amount of distributed data available on the user’s device to improve embeddings generation, while preserving user privacy. This approach will enhance the accuracy of semantic matching, reducing both false misses and false hits.
3.2. MeanCache’s System Design

### v2

@@ Abstract.
Large Language Models (LLMs) like ChatGPT and Llama2 have revolutionized natural language processing and search engine dynamics. However, these models incur exceptionally high computational costs. For instance, GPT-3 consists of 175 billion parameters where inference demands billions of floating-point operations. Caching is a natural solution to reduce LLM inference costs on repeated queries which constitute about 31% of the total queries. However, existing caching methods are incapable of finding semantic similarities among LLM queries, leading to unacceptable false hit-and-miss rates.
This paper introduces MeanCache, a user-centric semantic cache for LLMs that identifies semantically similar queries to determine cache hit or miss. Using MeanCache, the response to a user’s semantically similar query can be retrieved from a local cache rather than re-querying the LLM, thus reducing costs, service provider load, and environmental impact. Existing caching solutions for LLMs raise privacy and scalability concerns and perform wasteful query requests. MeanCache leverages Federated Learning (FL) to collaboratively train a query similarity model across LLM users without violating privacy. By placing a local cache in each user’s device and using FL, MeanCache reduces the latency and costs and enhances model performance, resulting in lower false hit rates. MeanCache compresses the embedding dimensions to minimize cache storage and also finds the optimal cosine similarity threshold. Our experiments benchmarked against the state-of-the-art caching method, reveal that MeanCache attains an approximately 17% higher F-score and a 20% increase in precision during semantic cache hit-and-miss decisions. It also reduces the storage requirement by 83% and accelerates semantic cache hit-and-miss decisions by 11%.
@@ 1. Introduction
Large Language Models (LLMs) like ChatGPT (cha, [n.d.]), Google Bard (bar, [n.d.]), Claude (cla, [n.d.]), and Llama 2 (Touvron et al., 2023) have demonstrated remarkable capabilities in understanding and generating human language, leading to significant advancements in applications ranging from search engines to conversational agents. LLMs are increasingly integrated into platforms like the Perplexity AI search engine  (per, [n.d.]b), Rabbit OS (rab, [n.d.]), Arc browser (arc, [n.d.]), and Bing Chat (bin, [n.d.]). For instance, in 2023, 1.9 billion queries were sent to Bing chat, reflecting the massive use of such systems (bin, [n.d.]).
@@
Motivation. Generating responses to user queries with LLMs, such as GPT-3, requires substantial computations and poses environmental challenges (Frantar et al., 2023; Wang et al., 2023; Tseng et al., 2024; Sachdeva et al., 2024; Ma et al., 2024; Strubell et al., 2019). For example, GPT-3’s 175B parameters in float16 format consume 326 GB memory, exceeding single GPU capacities and necessitating multi-GPU deployments (Frantar et al., 2023). These requirements lead to high operational costs.
Consequently, LLM-based services charge users and limit query rates (per, [n.d.]a; ope, [n.d.]). Prior studies report that users frequently submit similar queries to web services (Lempel and Moran, 2003; Xie and O’Hallaron, 2002)—approximately 33% of search engine queries being resubmitted (Markatos, 2001). We investigate if a similar pattern exists with queries to LLMs. Our analysis of a large corpus of real-world user conversations with ChatGPT reveals that approximately 31% of such queries have a high similarity with a previously submitted query 2. Thus, there is a potential to reduce the query inference cost of LLMs by up to one-third.
@@
Caching serves as an effective technique in traditional web services to address duplicate search queries, avoiding redundant computations, significantly improving response time, reducing the load on query processors, and enhancing network bandwidth utilization (Markatos, 2001; Saraiva et al., 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003; Brin and Page, 1998; Podlipnig and Böszörmenyi, 2003; Baeza-Yates et al., 2007). If applicable to LLMs-based web services, such caching can substantially impact billions of floating point operations, thereby decreasing operational costs and environmental impacts.
Figure 1. Analysis of real-world ChatGPT conversations reveals that, on average, 31% of queries are similar to previously submitted queries.
@@
Problem. Existing caching techniques  (Saraiva et al., 2001; Markatos, 2001; Xie and O’Hallaron, 2002; Lempel and Moran, 2003) use keyword matching, which often struggles to capture the semantic similarity among similar queries to LLM-based web services, resulting in a significantly low hit rate. For instance, existing caches do not detect the semantic similarity between ”How can I increase the battery life of my smartphone?” and ”Tips for extending the duration of my phone’s power source.”, leading to a cache miss. Recently, Zhu et al. (2023) and GPTCache(Bang, 2023) present server-side semantic caching for the LLMs-based services to address the limitations of keyword-matching caching techniques. If a new query is semantically similar to any query in the cache, the server returns the response from the cache. Otherwise, a model multiplexer selects the most suitable LLM for the query to generate the response.
@@
Existing semantic caches have several limitations. First, they demand a significantly large central cache to store the queries and responses of all users, which is unscalable and violates users’ privacy. Second, they incur the network cost of sending a user query to the server even if there is a cache hit. An end user will still be charged for the query even if the query is served from the server cache. Third, they use a single server-side embedding model to find the semantic similarity among queries, which does not generalize to each user’s querying patterns For instance, Google Keyboard (gbo, [n.d.]) adapts to each user’s unique writing style and embeds such personalized behaviors to enhance the accuracy of the next word prediction model. Fourth, they employ Llama-2 to enhance the accuracy of semantic matching (Bang, 2023); however, in practice, such models perform billions of operations to generate embeddings, offsetting the benefits of the cache.
@@
Key Insights and Contributions of MeanCache. This work introduces a novel user-centric semantic caching system called MeanCache. MeanCache provides a privacy-preserving caching system that returns the response to similar queries directly from the user’s local cache, bypassing the need to query the LLM-based web service. MeanCache achieves these goals in the following ways.
To address privacy concerns associated with central server-side caching, MeanCache introduces a user-side cache design ensuring that the user’s queries and responses are never stored outside of the user’s device. While a personalized per-user cache may not offer benefits for similar queries across multiple users, our investigation with real-world conversations with ChatGPT reveals that approximately 31% of the queries are still similar to at least one of the previously submitted queries by the same user. Therefore, having a user-centric cache offers the same, if not more, opportunities to eliminate expensive LLM queries.
To find a semantic match between a new query and cached queries, prior work (Bang, 2023) advocates for a larger model, such as Llama 2, to generate highly accurate embeddings but requires billions of operations. Such models are not suitable for resource-limited users’ devices. Previous work has shown that a smaller model can achieve performance comparable to larger models on custom tasks  (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). MeanCache uses smaller embedding models such as MPNet (Song et al., 2020) to generate embeddings for semantic matching locally.
Each user may not have sufficient queries to customize an embedding model that can help find a semantic match between new queries and cached queries. To address this, MeanCache utilizes federated learning, which exploits data silos on user devices for private training for collaborative learning, thereby personalizing an embedding model for each user. This privacy-preserving training not only customizes the embedding model to the user’s querying patterns but also enhances the performance (i.e., accuracy) of semantic caching without compromising user privacy (i.e., without storing user data on the web server).
The runtime performance of MeanCache is primarily influenced by the time taken to match a new query embedding vector with existing ones in the cache to find a semantically similar query. The search time is directly proportional to the dimensions of the embedding vector. To optimize runtime performance, MeanCache compresses the embedding vector by leveraging principal component analysis (PCA) (Pearson, 1901; Hotelling, 1933; Jolliffe and Cadima, 2016), effectively reducing the size of the embedding vector (i.e., projecting it to lower dimensional space). MeanCache also offers an adaptive cosine similarity threshold, which is also collaboratively computed using FL, to improve accuracy in finding semantic matches between queries.
Study of Real-World Queries. To measure the prevalence of similar queries in real-world conversation with LLM-based web service, We recruited 20 participants who regularly use ChatGPT and conducted a thorough analysis of their queries in a privacy-preserving manner. Across 27K queries, we find that 31% of a user’s queries are similar to at least one previous query by the same user, which could have been retrieved from a cache.
Evaluations. We compare MeanCache with GPTCache(Bang, 2023), a widely-used open-source semantic cache for LLM-based web services. GPTCache (Bang, 2023) is the most closely related work to MeanCache and has received over 6,000 stars on GitHub(zil, [n.d.]). We benchmark MeanCache’s performance against GPTCache on the GPTCache’s dataset (dat, [n.d.]) to demonstrate its effectiveness and highlight the improvements MeanCache offers over existing solutions.
MeanCache surpasses GPTCache (Bang, 2023) by achieving a 17% higher F score and approximately a 20% increase in precision in end-to-end deployment for identifying duplicate queries to LLM-based web services. MeanCache’s embedding compression utility approximately reduces storage and memory needs by 83% and results in 11% faster semantic matching, while still outperforming the state-of-the-art GPTCache.
2. Similar User Queries in Practice
Prior study (Markatos, 2001) measured the prevalence of similar queries across the entire user population. To understand the extent to which a single user submits similar queries, we conducted the first empirical study to measure how often a real-world ChatGPT user submits similar queries. Given the private nature of conversations with the ChatGPT service, study participants are naturally reluctant to share their query corpus. To address this problem, we develop a script that runs locally on each participant’s computer and identifies semantically similar queries, and calculates the percentage of similar queries on the participant’s query corpus. We then ask the participants to review the results and share, if comfortable, only three metrics (of similar queries, total queries, and days since the first query) anonymously with the authors.
We recruited 20 participants, including male and female participants, university professors, and graduate students who regularly use ChatGPT. Before measuring the query statistics, our analysis script addresses false positives by excluding large queries with lengths beyond 256 tokens. We examined over 27K queries across 20 participants who have been using ChatGPT for an average of 332 days.
Figure 1 presents the results of this study. We find that, on average, 31% of each participant’s queries are highly similar to at least one of the previously submitted queries by the same participant, which necessitates a user-centric cache for LLM-based web services.
@@ 3.1. Federated Learning
Federated Learning is a distributed, privacy-preserving training of a machine learning (ML) model. An FL round starts when a central server shares a global model with all the participating clients. Each client trains the model on its local data and sends the updated model to the central server. The central server aggregates the received models’ weights to form a new global model to be used in the next round. Different aggregating algorithms have been proposed, such as FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), and FedMA (Wang et al., 2020b). The FedAvg algorithm is the most popular one, and it aggregates the models with the following equation:
@@
$W_{global}^{t+1}$represents the new global model, $w_{k,t}$ denotes the model of the $k^{th}$ client at the $t^{th}$ round, $n_{k}$ are the number of samples available at $k^{th}$ client and $n$ is the total number of samples of clients participating in the given round. This procedure is reiterated for multiple rounds until the model reaches convergence.
@@ 3.2. Transformer Architecture
The transformer architecture is a deep learning model based on the attention mechanism, which allows the model to focus on the important parts of the input (Vaswani et al., 2017). The transformer architecture is primarily used for sequence-to-sequence tasks in NLP, such as machine translation, text summarization, and question-answering.
The attention mechanism calculates the importance of each input element by considering its relationship with other elements in the sequence. This enables the model to capture long-range dependencies and improve performance on tasks that require an understanding of context. The feed-forward layer
helps the model capture complex patterns and relationships in the data. The transformer architecture also incorporates positional encoding, which provides information about the position of each element in the input sequence. This helps the model understand the order of the elements, which is crucial for tasks like machine translation.
@@ 3.2.1. Embeddings and Cosine Similarity
Figure 2. MeanCache’s Workflow
4. MeanCache’s Design
MeanCache is a user-centric semantic caching system optimized for user-side operation. Figure 2 illustrates the workflow of MeanCache for similar queries. When a user submits a query to an LLM web service with MeanCache enabled, MeanCache computes the query embeddings. These embeddings are then matched with the embedding of the cached queries using cosine similarity. If MeanCache locates a similar query within the cache, the response is retrieved from the local cache and returned to the user. Otherwise, MeanCache forwards the query to the LLM web service to obtain the response. The query and its response and embeddings are then stored in the cache.